OpenStudy (anonymous):

limit as x approaches 0 of ((2-h)^5-32)/(h)

OpenStudy (anonymous):

Just wondering - do you know how to compute derivatives? Or are you just learning how to do that and have to rely on the limit definition?

OpenStudy (anonymous):

I know derivatives

OpenStudy (anonymous):

I like derivatives better. :) haha

OpenStudy (anonymous):

Ok, that makes things easier then. Note that \(\large \displaystyle\lim_{h\to 0}\frac{(2-h)^5-32}{h}\) looks something like \(\large\displaystyle\lim_{h\to 0}\frac{f(a+h)-f(a)}{h}=f^{\prime}(a)\); in particular, \[\large\lim_{h\to 0}\frac{(2-h)^5-32}{h} = -\lim_{h\to 0}\frac{((-2)+h)^5 -(-32)}{h}=\ldots\]Can you take things from here? :-)

OpenStudy (anonymous):

wait so I can just find the derivative?

OpenStudy (anonymous):

Yup, that's the trick here. XD But the derivative of what function at what point?

OpenStudy (anonymous):

Hmm I am not sure

OpenStudy (anonymous):

So I take the derivative and then evaluate that limit at 0

OpenStudy (anonymous):

Would it be clearer if I told you that \[\large \lim_{h\to 0}\frac{(2-h)^5-32}{h}= \lim_{h\to 0}\frac{[-((-2)+h)^5] - [-(-2)^5]}{h}?\]

OpenStudy (anonymous):

so then i evaluate it as h approaches zero. Or do I not do that?

OpenStudy (anonymous):

The point of this "trick" is to avoid computing a limit completely and instead evaluate a derivative at a point.

OpenStudy (anonymous):

The observation to be made here is that the limit your given looks like the difference quotient of some function at a certain point. The objective then is to determine what function you're differentiating and at what point.

OpenStudy (anonymous):

There's nothing wrong with going ahead and expanding everything out and then simplify the whole thing; it's just that this method saves you a lot of time, especially if you're familiar with derivatives at this point.

OpenStudy (anonymous):